Suppose lim n→∞ (sn −1)/(sn +1)=0. Prove that limn→∞sn = 1


1
Expert's answer
2021-02-03T05:02:21-0500

Here we have,


"\lim\limits_{x\to\infin}\frac{s_n-1}{s_n+1}=0\\~\\
\Rightarrow \lim\limits_{x\to\infin} \frac{s_n+1-2}{s_n+1}=0\\~\\
\Rightarrow \lim\limits_{x\to\infin}(1-\frac{2}{s_n+1})=0\\~\\
\Rightarrow 1-\lim\limits_{x\to\infin}\frac{2}{s_n+1}=0\\~\\
\Rightarrow \lim\limits_{x\to\infin}\frac{1}{s_n+1}=\frac{1}{2}"

Now, as this is an equality and RHS is finite, we can be sure that "\lim\limits_{x\to\infin}s_n" is some value let say "c\in\R" .

i.e. "c=\lim\limits_{x\to\infin}s_n"


Therefore,



"\frac{1}{c+1}=\frac{1}{2}\\~\\
\Rightarrow c+1=2\\~\\
\Rightarrow c=1"

So, we have:-



"\fcolorbox{black}{aqua}{$\textcolor{black}{\lim\limits_{x\to\infin}s_n=1}$}"